# Intelligent Tutors for Adult Learners: An Analysis of Needs and Challenges

### v2

@@ Abstract.
This research examines the sociotechnical factors that influence the adoption and usage of intelligent tutoring systems in self-directed learning contexts, focusing specifically on adult learners. The study is divided into two parts. First, we present Apprentice Tutors, a novel intelligent tutoring system designed to address the unique needs of adult learners. The platform includes adaptive problem selection, real-time feedback, and visual dashboards to support learning in college algebra topics. Second, we investigate the specific needs and experiences of adult users through a deployment study and a series of focus groups. Using thematic analysis, we identify key challenges and opportunities for improving tutor design and adoption. Based on these findings, we offer actionable design recommendations to help developers create intelligent tutoring systems that better align with the motivations and learning preferences of adult learners. This work contributes to the broader understanding of how to enhance educational technologies to support lifelong learning and professional development.
@@ 1. Introduction
Intelligent tutoring platforms have come a long way over the past five decades, since the earliest mentions of tutors in the 1970s. Intelligent tutors support users by guiding and scaffolding problem-solving across a wide variety of problems and generalized problem types (MacLellan and Koedinger, 2022). One of the earliest intelligent tutor based on an expert model was GUIDON, developed by William Clancey to teach medical knowledge (Clancey, 1986). Since then, research on intelligent tutors has expanded, with K-12 education becoming a central focus, particularly in environments where learners are tasked with using tutors as part of their curriculum (Koedinger et al., 1997). Several large randomized controlled trials have demonstrated the effectiveness of tutors in diverse populations, topics, and contexts (Pane et al., 2014b; VanLEHN, 2011; Wijekumar et al., 2012).
Despite a consistent track record of improving student learning, tutors have not been widely adopted among adult learners across non-traditional educational environments, such as workplace training programs, online education platforms, and continuing professional development courses. Adult learners are typically defined as non-traditional students over the age of 18 who engage in continued education beyond undergraduate studies. These learners often juggle multiple responsibilities, such as full-time employment or family commitments. Recent research on the adoption of educational technologies by adult learners indicates that this population is still significantly underexplored (Gasevic et al., 2021). In specialized contexts, however, there has been work exploring the use of tutoring systems to support adult learners. For example, the Generative Instructional Framework for Training (GIFT) framework was developed to tailor tutors to the unique needs of military personnel, focusing on skills acquisition and decision making under pressure (Sinatra et al., 2014). Unlike K-12 students, who are often required to use tutors as part of their curriculum, adult learners are driven by intrinsic motivations. This highlights the importance of investigating how tutors are utilized in voluntary, non-mandated settings, where engagement stems from personal goals rather than external obligations.
@@
One line of research proposes that tutors have not been widely adopted due to the high cost of creating them (MacLellan and Koedinger, 2022; Murray et al., 2003). While we acknowledge that tutor authoring challenges contribute to a lack of adoption, this research explores an additional hypothesis: that tutors are often not adopted due to sociotechnical challenges that arise from a mismatch between the perceptions and needs of stakeholders—mainly learners—and the key design features and capabilities of intelligent tutors. As such, there were two research questions that inspired the work in this field:
  1. (1)
    RQ1: What are adult learners intelligent tutoring needs?
  2. (2)
    RQ2: How might we craft intelligent tutors that better align with adult learner needs?
To answer these research questions, this study is split into two parts. The first part presents a novel intelligent tutoring system, Apprentice Tutors, which is designed for the adult learner population. The second part concentrates on the deployment of the Apprentice Tutors to a population of adult learners at a state technical college. This study specifically concentrates on tutor adoption by a specific learner population. Adult learners are the main focus of this study, as their motivations and learning approaches differ significantly from those of K-12 students. Although younger learners typically engage in structured curriculum and are driven by external factors such as grades or parental encouragement, adult learners are often self-directed in their educational pursuits. This self-directedness involves taking the initiative to identify learning needs, set personal and professional goals, and seek resources aligned with these objectives (Kanfer et al., 2022). Recent research highlights the critical role of self-management in adult learning, emphasizing that the ability to regulate learning activities is essential to acquire new skills and advance career development (Kanfer et al., 2022).
Furthermore, previous studies, including large randomized trials, focused mainly on K-12 learners, necessitating investigation into adult use of pedagogical tools on a scale (Koedinger et al., 1997; Heffernan and Heffernan, 2014; Aleven et al., 2006). In particular, previous studies on tutor adoption have limited their focus on measuring tutor effectiveness, without confirming whether tutor effectiveness equates to tutor adoption (VanLehn et al., 2005; Bowen et al., 2012). In fact, the influence of other factors on the adoption of tutors by adult learners is largely unexplored.
The Apprentice Tutors platform was created to address the first research question. It covers four mathematical topics: radicals, exponents, factoring polynomials, and rational equations. All tutors offer a learning experience that includes many features typical of intelligent tutoring systems, including on-demand hints, real-time correctness feedback, adaptive problem selection, and a multilayer hinting strategy. Later in this paper, the inner workings of the Apprentice Tutors platform are explored in greater detail. The tutoring materials were developed based on the existing course curriculum and refined based on iterative feedback from the course instructors, who also provided anecdotal evidence about the areas where the students faced difficulties. Furthermore, in the first study presented, these tutors were deployed to adult users, who had the option to use any of the Apprentice Tutors during their College Algebra course. The tutors were deployed to the learners via their learning management system (Blackboard) over the course of an academic year, divided into four quarters: fall, winter, spring, and summer. Tutors were embedded into students’ course curriculum, but were not mandatory. In this context, adult learners had more flexibility in whether they wanted to use the tutors. Through the deployment, the frequency of tutor adoption was analyzed.
@@
The second study aims to better understand how these tutors were perceived by end users. To investigate the unique needs of users, focus groups were conducted with both students and teachers. Following the focus groups, a group of five graduate students worked together to transcribe the focus group sessions and conducted qualitative analysis of these data to derive key themes that emerged around the specific challenges and needs faced in adult learning contexts. In addition, thematic analyzes were performed on the data collected from the previous focus group (Braun and Clarke, 2006). To this end, key themes were synthesized into generalized recommendations on how to improve the usage and adoption of intelligent tutors in adult learning contexts. The aim of this work is to guide designers and developers in building intelligent tutors to better understand the needs of end-users and build tools that are susceptible to greater overall adoption and impact.
@@ 2. Related Works
Intelligent tutors are praised for their scale-based personalized teaching abilities that address limitations that often hinder access to individualized high-quality tutoring. Although personalized human tutoring is particularly effective, it is cost prohibitive and therefore not available everywhere. Some researchers argue that private tutoring can disproportionately benefit those with the financial means to afford such services, thereby exacerbating existing educational inequities (Bray and Lykins, 2012; Bray, 2013; Tansel and Bircan, 2006; Byun and Park, 2012). There have been several studies where intelligent tutors have shown success when used by K-12 learners in controlled settings where all participants engage in a fixed amount of activity (Koedinger et al., 1997; Heffernan and Heffernan, 2014; Aleven et al., 2006; Razzaq et al., 2009; Pardos et al., 2023).
@@
For example, a randomized trail with an algebra tutor deployed as part of the Pittsburgh Urban Math Project. It found that, on average, 470 students in the tutored group outperformed students in the non-tutored group by 15% on standardized tests and achieved an improvement 100% in concepts that the tutoring system targeted (Koedinger et al., 1997). Similarly, a large-scale randomized controlled trial of Carnegie Learning’s Cognitive Tutor Algebra I found significant improvements in student achievement, with tutored students outperforming their peers in traditional classrooms, specifically in middle and high school math (Pane et al., 2014a).
@@
While certain pedagogical principles remain consistent across age groups, such as the benefits of high-quality feedback (Hattie and Timperley, 2007; Shute, 2008; Nattamai Subramanian Rajkumar et al., 2024), it is important to recognize that adult learners differ from K-12 learners in ways that significantly impact the learning process. The context of adult learning often diverges from traditional instructional settings, frequently being self-directed rather than formal (Kanfer et al., 2022). From a cognitive development perspective, some abilities, such as memory and abstract reasoning, tend to decline after early adulthood, while others, such as crystallized knowledge, continue to increase through midlife (Baltes, 1997; Schaie, 2005). These developmental changes require considerations for age-inclusive instruction. Considering both the cognitive changes of the lifespan and the contextual differences in learning environments, it becomes evident that the research and practices developed for tutors may not be fully translated to adult learners. Although organizational researchers have provided guidelines for instructional design in technology-supported training contexts (Salas et al., 2012), it remains unclear how or to what extent these principles have been applied to develop effective tutors for adults. This knowledge gap highlights the need for further research into the specific requirements and considerations for designing tutors that meet the unique needs and characteristics of adult learners.
This study addresses two key gaps in the research of intelligent tutoring systems for adult learners: (1) limited information on why adult learners choose to use tutors and (2) insufficient design recommendations for self-directed adult learners. To address the first gap, we developed the Apprentice Tutors deployed them within College Algebra classes at a state technical college to investigate factors influencing voluntary adoption and usage. To address the second gap, we analyzed focus group data to derived design recommendations based on adult learners’ needs in self-directed contexts. This work contributes empirically grounded insights into the participation of adult learners in tutoring and offers guidelines to improve intelligent tutor design and adoption in adult learning environments, thus addressing a notable gap in the existing literature (Shi et al., 2021). The findings aim to guide the development of intelligent tutors that better align with the expectations and requirements of adult learners in voluntary educational settings.
@@
3. Apprentice Tutors
3.1. Platform Description
@@
The Apprentice Tutors platform, presented in Figure 1, is a web-based intelligent tutoring system for adult learners that is accessible via most modern browsers from a mobile device or a computer. Each tutor within the platform consists of multiple problem types, where each problem type consists of an expert model, which can provide feedback on student input, and an interface to scaffold problem solving. The Apprentice Tutors architecture was built with Python and the expert models were built with a rete-based production rule engine (Forgy, 1989).
@@
To develop tutors on this platform, a cognitive task analysis was performed with instructors (Lovett, 1998). Based on the information from this analysis, the tutor interface was designed and programmed for each specific problem type. Each step within the tutor interface is mapped to a single skill, which is represented in the expert model via a particular production rule.
@@
As users interact with steps in a tutor interface, each tutor transaction is stored within the Apprentice Tutor platform’s database. This allows it (the tutoring platform) to track the progression of skills over time as users continue to interact (Vanlehn, 2006). We created an adaptive problem selection feature to personalize the learning experience (Vanlehn, 2006). Depending on how they would like to practice, learners can choose between adaptive problem selection (specified in Section 1.3.6) or manually select a particular problem type within each tutor. Each tutor also generates randomized problems for students to practice based on the user interface constraints.
@@
As users interact with dynamically generated problems, the tutors provide real-time correctness feedback. The tutors also provide a series of multi-layer hints (specified in Section 1.3.5) in a text box below the tutor interface. The Apprentice Tutors platform also has visualization dashboards that allow both students and teachers to track skill mastery (specified in Section 1.3.7) (Guo et al., 2024). Apprentice Tutors allows for the seamless deployment of tutors to classrooms. To facilitate this goal, the interoperability standards for learning tools developed by the IMS Global Learning Consortium (Lynch and McLean, 2004) were used. All major learning management systems comply with these standards and allow educational technologies to be embedded in popular learning management systems, such as Blackboard and Canvas.
3.2. Apprentice Tutors Topics
@@
In this study, we developed four tutors for four college algebra topics: radicals, exponents, factoring polynomials, and rational equations (see Figure 1). The course uses the OpenStax college algebra textbook and the tutors are designed to support topics from this book (OpenStax, 2020). I created the initial problems in the Apprentice Tutors to align with the worked examples in the course textbook. For example, in the exponent unit, the problem sets included the exponent power rule, the exponent product rule, and the exponent quotient rule, leading to the creation of individual tutor problems for each of these problem types.
3.3. Building Expert Models
A tutoring system that incorporates an expert model is referred to as a cognitive tutor (Anderson et al., 1995). An expert model is capable of generating multiple solutions to the problems presented to students, enabling personalized and adaptive feedback. An expert model is a core component of cognitive tutors, representing the idealized knowledge and problem solving strategies of a domain expert. The expert model serves as a benchmark against which a learner’s performance can be evaluated, enabling the tutor to provide real-time feedback and hints. In combination with a knowledge tracing algorithm (Corbett and Anderson, 1994), the expert model can also support adaptive problem selection. Expert models are constructed using formalized rules or algorithms that encapsulate the domain knowledge required to solve problems within the tutor.
Constructing an expert model for a complex tutor interface with many steps can be challenging. The development process is further complicated by interdependencies among steps in the problem scaffolding, where solving one step correctly depends on prior steps. Another difficulty arises when creating generalized expert models that operate effectively in both unrestricted and problem-specific scenarios (Anderson et al., 1995).
In the Apprentice Tutors, each production rule represents a unit of knowledge or a specific skill required to complete a step in the problem. I authored production rules for each step of each tutor, and the production engine (a Python-based system implementing the algorithm outlined by (Doorenbos, 1995)) dynamically composes these rules at runtime to accommodate the full range of variation in inputs. Cognitive tutors have been shown to enhance learning outcomes by providing immediate, personalized feedback, and adaptive problem sequencing, making them highly effective for individualized instruction. Despite the inherent challenges of constructing an expert model, I adopted this approach to power the tutors, ensuring they could model the procedural knowledge with precision and flexibility (Anderson et al., 1995; Vanlehn, 2006; Koedinger et al., 1997).
Each problem type within a tutor is associated with a corresponding problem interface, expert model, and problem generator. The expert model enables the tutor to provide correctness feedback, guiding learners by evaluating their input against the model rules. It also supports the generation of context-sensitive hints, offering step-by-step guidance when learners encounter difficulties. Although programming an expert model for each type of problem is time-intensive, this approach ensures that tutors can function on a wide variety of randomly generated problems, enhancing their robustness and adaptability.
By establishing the expert model in the principles of cognitive theory and using proven techniques from the intelligent tutoring literature, this system design ensures both theoretical soundness and practical effectiveness. These choices reflect a broader commitment to creating tutors that align with established frameworks while addressing the unique challenges of complex problem domains.
3.4. Real-Time Correctness Feedback
Real-time feedback has been shown to be an effective instructional strategy, particularly in improving learning outcomes and fostering immediate error correction (Shute, 2008). By providing learners with timely feedback, the tutor helps them identify and correct errors before they become ingrained, promoting the mastery of the targeted skills. Research in intelligent tutoring systems has consistently shown that immediate feedback improves engagement and supports learners to develop a deeper understanding of the material (Koedinger and Anderson, 1997).
@@ 3.4. Real-Time Correctness Feedback
The production rules in the system form the foundation of this feedback mechanism and have two key components: conditions and functions. When a learner enters a value into a tutor interface, the system analyzes the tutor’s state prior to their input, along with the input itself. All productions with conditions that match the current tutor state are collected, and their associated functions are executed. The system supports chaining multiple rules together to interpret and explain student input. For instance, each rule is designed to correspond to a specific step in the interface, with computations performed by the rules stored as variables. These stored variables can then be referenced in subsequent production rules, eliminating the need to re-implement the same computations. If the student’s input matches the output of the specific production function linked to the interface, the input is marked as correct. In this case, the corresponding field is highlighted in green and is disabled to prevent further editing, prompting the user to proceed to the next step. Conversely, if the student’s input does not match the output of any production function, it is marked as incorrect, and the field is highlighted in red, prompting the user to retry the step. This functionality enables the tutor to provide real-time correctness feedback to users (as shown in Figure 1, part B). Using production rules to deliver immediate and actionable feedback, Apprentice Tutors ensure that learners receive guidance tailored to their specific input.
3.5. Hinting Strategy

### v3

@@ Abstract.
This work examines the sociotechnical factors that influence the adoption and usage of intelligent tutoring systems in self-directed learning contexts, focusing specifically on adult learners. The study is divided into two parts. First, we present Apprentice Tutors, a novel intelligent tutoring system designed to address the unique needs of adult learners. The platform includes adaptive problem selection, real-time feedback, and visual dashboards to support learning in college algebra topics. Second, we investigate the specific needs and experiences of adult users through a deployment study and a series of focus groups. Using thematic analysis, we identify key challenges and opportunities to improve tutor design and adoption. Based on these findings, we offer actionable design recommendations to help developers create intelligent tutoring systems that better align with the motivations and learning preferences of adult learners. This work contributes to a wider understanding of how to improve educational technologies to support lifelong learning and professional development.
@@ 1. Introduction
Intelligent tutoring systems support learners by guiding and scaffolding problem solving across a variety of domains (MacLellan and Koedinger, 2022). Tutors have a well-documented history of improving student learning, particularly in K-12 settings (Pane et al., 2014), given to their ability to provide personalized instruction, immediate feedback, and adaptive support. In theory, these benefits should extend to adult learners in non-traditional educational contexts — such as workplace training, online education platforms, and continuing professional development. However, tutors remain largely underutilized in these settings, either due to a lack of adoption or a limited body of research on their effectiveness for adult learners (Bernacki et al., 2021). Our work seeks to address this gap by exploring how tutors can support learning in adult education.
While the term adult learner colloquially refers to students over the age of 18, we define it here as individuals who are both over 18 and classified as non-traditional students—those pursuing education outside the conventional schooling pathway (e.g., a four-year undergraduate degree immediately after high school). This definition includes students who enter the workforce after high school and then return to college to change careers (e.g., through a two-year technical degree). Unlike traditional students, adult learners often juggle multiple responsibilities, such as full-time employment or family commitments. Recent research on the adoption of educational technologies by adult learners indicates that this population remains understudied (Bernacki et al., 2021).
Unlike K-12 students, who are often required to use tutors as part of their curriculum, adult learners self-regulate their own learning and are more strongly driven by intrinsic motivation (Lyndgaard et al., 2024). This highlights the importance of investigating how tutors are utilized in voluntary, non-mandated settings, where engagement stems from personal goals rather than external obligations.
@@
While prior work on tutor adoption across a variety of user segments hypothesizes that tutors have not been widely adopted due to their high cost of development (MacLellan and Koedinger, 2022; Murray et al., 2003), this work explores an alternative hypothesis: that tutors are not adopted due to sociotechnical challenges that arise from a mismatch between the perceptions and needs of learners and the capabilities of intelligent tutors. Accordingly, our work is guided by two research questions:
  1. RQ1:
    Will adult learners use intelligent tutors when provided as a supplementary learning resource?
  2. RQ2:
    What are the specific needs of adult learners when using educational tools?
@@ 2. RQ2:
To explore both questions, we developed an intelligent tutoring system based on best practices, called Apprentice Tutors. We then conducted a two-part study. The first part introduces Apprentice Tutors and examines whether students adopt the system. The second part evaluates whether the tutor meets users’ needs and explores what insights can be gained about adult learning needs from its deployment. To investigate the unique needs of users, we conducted focus groups with both students and teachers. Following the focus groups, we performed a thematic analysis of the data (Braun and Clarke, 2006) and synthesized key findings into recommendations to improve the adoption and effectiveness of intelligent tutors in adult learning. The aim of this work is to guide designers and developers in building intelligent tutors with a better understand of needs of end-users and building pedagogical tools that allow for greater overall adoption and impact.
@@ 2. Related Works
Intelligent tutors are praised for their ability to support personalized learning at scale. Although human tutoring is particularly effective, it is cost prohibitive and therefore not available to everyone. Some researchers argue that private human tutoring can disproportionately benefit those with the financial means to afford such services, thereby exacerbating existing educational inequities (Bray and Lykins, 2012; Bray, 2013; Tansel and Bircan, 2006; Byun and Park, 2012). There have been several studies where intelligent tutors have shown success when used by K-12 learners in controlled settings where all participants engage in a fixed amount of activity (Koedinger et al., 1997; Heffernan and Heffernan, 2014; Aleven et al., 2006; Razzaq et al., 2009; Pardos et al., 2023).
@@
For example, a randomized trial involving an algebra tutor was conducted as part of the Pittsburgh Urban Math Project. The study found that, on average, 470 students in the tutored group outperformed students in the non-tutored group by 15% on standardized tests and achieved a 100% improvement in the concepts targeted by the tutoring system (Koedinger et al., 1997). Similarly, a large-scale randomized controlled trial of Carnegie Learning’s Cognitive Tutor Algebra I found significant gains in student achievement, with tutored students outperforming their peers in traditional classrooms (Pane et al., 2014).
@@
While certain pedagogical principles remain consistent across age groups, such as the benefits of high-quality feedback (Hattie and Timperley, 2007; Shute, 2008; Nattamai Subramanian Rajkumar et al., 2024), it is important to recognize that adult learners differ from K-12 learners in ways that affect learning. The context of adult learners diverges substantially. Often, they have to balance education with many other life factors. Further, adult learners are much more self-directed and intrinsically motivated than K-12 learners (Kanfer et al., 2022). From a cognitive development perspective, some abilities, such as memory and abstract reasoning, tend to decline after early adulthood, while others, such as crystallized knowledge, continue to increase through midlife (Baltes, 1997; Schaie, 2005). These developmental changes require considerations for age-inclusive instruction. Considering both the cognitive changes of the lifespan and the contextual differences in learning environments, it becomes evident that the research and practices developed for tutors may not be fully transferrable to adult learners. Although organizational researchers have provided guidelines for instructional design in technology-supported training contexts (Salas et al., 2012), it remains unclear how or to what extent these principles have been applied to develop effective tutors for adults. This knowledge gap highlights the need for further research into the specific requirements and considerations for designing tutors that meet the unique needs and characteristics of adult learners.
@@
3. Study 1: Building and Deploying Apprentice Tutors
Much of existing research on educational technology focuses on deploying intelligent tutors in K–12 settings (Koedinger et al., 1997; Heffernan and Heffernan, 2014; Aleven et al., 2006). However, it remains unclear whether these tutors have the same impact on adult learners when used as a supplementary resource. To address this gap and our first research question, we developed an intelligent tutoring system specifically for adult learners, where participation is voluntary. By analyzing tutor adoption and usage over time, this study provides insights into how this understudied user group interacts with educational tools.
3.1. Apprentice Tutors
@@
We developed a web-based intelligent tutoring system platform, called Apprentice Tutors, that is accessible via most modern browsers from a mobile device or a computer, see Figure 1. Each tutor supports tutoring of multiple types of problems. Each problem type consists of an interface to scaffold problem solving and an expert model that can provide feedback on student input. The Apprentice Tutors architecture was built with Python and the expert models were built with a rete-based production rule engine (Forgy, 1989).
@@
To develop tutors on this platform, we performed a cognitive task analysis with instructors (Lovett, 1998). Based on the information from this analysis, we designed and programmed a tutor interface for each specific problem type. Each step within the tutor interface is mapped to a single skill, which is represented in the expert model via a particular production rule.
@@
Each tutor transaction is stored within the Apprentice Tutor platform’s database. This allows it to track the progression of skills over time as users continue to interact. We created an adaptive problem selection feature to personalize the learning experience (Vanlehn, 2006). Depending on how they would like to practice, learners can choose between adaptive problem selection or manually select a particular problem type within each tutor. Each tutor also generates randomized problems for students to practice based on the user interface constraints.
@@
As users interact with dynamically generated problems, the tutors provide real-time correctness feedback and multi-layer hints in a text box below the tutor interface. The Apprentice Tutors platform also has visualization dashboards that allow both students and teachers to track skill mastery (Guo et al., 2024). Apprentice Tutors allows for the seamless deployment of tutors to classrooms. To facilitate this goal, the interoperability standards for learning tools developed by the IMS Global Learning Consortium (Lynch and McLean, 2004) were used. All major learning management systems comply with these standards and allow educational technologies to be embedded in popular learning management systems, such as Blackboard and Canvas.
3.1.1. Apprentice Tutors Topics
@@
For this study, we developed four tutors covering radicals, exponents, factoring polynomials, and rational equations (Figure 1). Aligned with the OpenStax college algebra textbook (OpenStax, 2020), the tutors support textbook topics. Problem sets were designed to match worked examples, such as creating individual tutor problems for exponent rules like power, product, and quotient.
3.1.2. Building Expert Models
A tutoring system incorporating a rule-based expert model is known as a cognitive tutor (Anderson et al., 1995). The expert model generates multiple solutions, enabling personalized, adaptive feedback by evaluating student performance against domain-expert strategies. Combined with a knowledge tracing algorithm (Corbett and Anderson, 1994), it also supports adaptive problem selection. Building expert models for complex tutors is challenging due to step interdependencies and the need for problem-specific and generalized reasoning (Anderson et al., 1995). In Apprentice Tutors, each production rule represents a skill required to complete a step. The production engine, based on (Doorenbos, 1995), dynamically composes these rules at runtime to handle varied inputs. Cognitive tutors improve learning by providing immediate, personalized feedback, and adaptive problem sequencing, making them effective for individualized instruction (Anderson et al., 1995; Vanlehn, 2006; Koedinger et al., 1997). Each type of problem in Apprentice Tutors has a dedicated interface, expert model, and problem generator. The expert model enables correctness feedback and context-sensitive hints. While developing expert models is time-intensive, it ensures tutors function across randomly generated problems, enhancing adaptability. Grounded in cognitive theory and intelligent tutoring research, this design balances theoretical soundness with practical effectiveness.
3.1.3. Real-Time Correctness Feedback
@@
Real-time feedback has been shown to be an effective instructional strategy, particularly in improving learning outcomes and fostering immediate error correction (Shute, 2008). By providing learners with timely feedback, the tutor helps them identify and correct errors before they become ingrained, promoting the mastery of the targeted skills. With the Apprentice Tutors, each rule is designed to correspond to a specific step in the interface, with computations performed by the rules stored as variables. If the student’s input matches the output of the specific production function linked to the interface, the input is marked as correct. In this case, the corresponding field is highlighted in green and is disabled to prevent further editing, prompting the user to proceed to the next step. Conversely, if the student’s input does not match the output of any production function, it is marked as incorrect, and the field is highlighted in red, prompting the user to retry the step. This functionality enables the tutor to provide real-time correctness feedback to users (as shown in Figure 1, part B).
3.1.4. Hinting Strategy